What is the volume of the solid obtained by rotating the region under the curve y=cosx between x=0 and x=π/12 about the x-axis.
1
Expert's answer
2020-06-22T18:18:53-0400

The volume of solid obtained by rotation of the curve "f(x)" between "x=a" and "x=b" about the x-axis can be calculated as

"V = \pi \int\limits_a^bf^2(x)\,dx" . In our case

"V = \pi \int\limits_0^{\pi/12}\cos^2 x\,dx = \pi \int\limits_{0}^{\pi/12}\dfrac{1+\cos(2x)}{2}\,dx = \pi \int\limits_{0}^{\pi/12}\dfrac12\,dx + \pi\int\limits_{0}^{\pi/12} \dfrac{\cos(2x)}{2}\,dx = \dfrac{\pi^2}{24} + \pi \int\limits_{0}^{\pi/6}\dfrac{\cos \tau}{4}\,d\tau = \dfrac{\pi^2}{24} + \dfrac{\pi}{4} \sin\tau\Big|_0^{\pi/6} = \dfrac{\pi^2}{24} +\dfrac{\pi}{8}."